Use the compund interest formulas A=P(1+r/n)^nt and A= Pe^rt to solve.
Find the accumulated value of an investment of $13,000 at 8% compounded semiannually for 5
years.
a. $15,816.49
b. $19,101.26
c. $18,200.00
d. $19,243.18

Answer by stanbon(75887)

Use the compund interest formulas A=P(1+r/n)^nt and A= Pe^rt to solve.
Find the accumulated value of an investment of $13,000 at 8% compounded semiannually for 5 years.
---------------------------
A(5) = 13000(1+(0.08/2))^(2*5)
A(5) = $19243.18
=====================
Cheers,
Stan H.
==============
a. $15,816.49
b. $19,101.26
c. $18,200.00
d. $19,243.18